HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

SPECTRA OF THE FORMALDEHYDE MOLECULE IN THE RANGE 3800−5100 CM^{−1}

ALENA RODINA, ANDREI V. NIKITIN, Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055 Russia; \underline{MICHAËL REY} , Groupe de Spectrométrie Moléculaire et Atmosphérique UMR CNRS 7331, UFR Sciences BP 1039, 51687 Reims Cedex 2, France; \underline{LAURENT MANCERON} , Université Paris-Est Créteil,

CNRS, LISA, F-75013 Paris, France.; ALEXANDER PROTASEVICH, \underline{VLADIMIR TYUTEREV} , Laboratory of Theoretical Spectroscopy, V.E. Zuev Institute of Atmospheric Optics SB RAS 1, Akademician Zuev Sq., Tomsk, 634055

Russia

A comprehensive study of vibration-rotation spectral patterns of the formaldehyde molecule will be reported. Vibrational energy levels of H_{2}CO were computed using variational nuclear motion calculations from ab initio and empirically optimized full 6-dimensional potential energy surfaces in the electronic ground state. Ab initio calculations were carried out using extended electronic structure coupledcluster method accounting for the dynamic electron correlations including triple and quadruple excitations as well as relativistic and diagonal Born-Oppenheimer corrections^{1}. Effective Hamiltonian and effective dipole transition moments for the rovibrational polyads were derived from the ab initio surfaces using high-order contact transformations (CT) ^{2}^{,}^{3}. The analyses of experimental spectra recorded at T=160−165 K and various pressures were performed for the region of the 3-rd vibrational polyad. The experimental setup was described in Ref.^{4}. The line positions and intensities were obtained by nonlinear least-squares fits with the SpectraPlot software^{5} and analyzed using the CT-models for the effective Hamiltonian and effective dipole transition moments. Fifteen new vibrational levels were derived from these analyses in a good agreement with calculations^{1}. In the range 4500−4700 cm^{−1}, a line list with the quantum identification was produced. It can be useful for a global modelling of the formaldehyde spectra and for an improvement of the line parameters in the databases^{6}^{,}^{7}^{,}^{8}.

1A.V. Nikitin, A.E. Protasevich, A.A. Rodina, M. Rey, A. Tajti, Vl.G. Tyuterev Y, J. Quant. Spectrosc. Radiat. Transfer 260, 107478 (2021).

2V. Tyuterev, S. Tashkun, M. Rey, A. Nikitin, Mol. Phys., 120, issue 15-16 art. no. e2096140 (2022)

3M. Rey, J. Chem. Phys, 292, 108349 (2022)

4F.K. Tchana, F. Willaert, X. Landsheere, J.-M. Flaud, L.Lago, M. Chapuis, C. Herbeaux, P. Roy, L. Manceron, Review Scientif. Instrum. 84, 093101 (2013).

5A. V. Nikitin, R. V. Kochanov, Optic. Atmos. 24, 931-941 (2011).

6I.E. Gordon et al., J. Quant. Spectrosc. Radiat. Transfer 277, art. no. 107949 (2022).

7T. Delahaye, et al., J. Mol. Spectrosc. 380, art. no. 111510 (2021).

8M. Rey, A. Nikitin, Y. Babikov, V. Tyuterev, J. Mol. Spectrosc, 327, 138-158 (2016)

p-number: p100  Submitted on Tue, 13 Jun 2023 15:31:12 +0200

This work was supported by the Russian Scientific Foundation (RSF, No. 22-42-09022) and Agence Nationale de la Recherche (ANR, French National Research Agency, Grant No. 21-CE30-0053-01).

p-number: p100  Submitted on Tue, 13 Jun 2023 15:31:12 +0200